# Finite Dynamical Laminations

### v1

@@ 5 What are the limit sets of FDL?
Of course, we wish to classify these sequences by how nice their limits are. For that purpose we summarize the definitions of a few terms which we use sparsely. The most important term is q-lamination, which is defined in [1] as a lamination having an equivalent relation that is laminational with the added convention that a leaf is in the lamination iff it is in the boundary of the convex hull of an equivalence class. Laminational, is also defined in [1], and refers to an equivalence relation that meets a number of conditions that are each necessary for the lamination to be realized by a polynomial in the sense contemplated here. One such condition is that the equivalence classes are finite. These terms generally presuppose that the lamination is invariant. Unclean is defined in Thurston as having 3 leaves meet at a point. Unclean laminations are not q, but they might be proper, in which case their equivalence relations are laminational. When we say that a lamination is not proper we mean that it has infinite equivalence classes. To remember this, think that an unclean sibling lamination can be cleaned into a q-lamination iff it is proper.
@@ 6 Are all FDL realized by a polynomial?
Kiwi’s theorem does not show that all rays land in the way described by the lamination but instead that the impression of each ray intersects as described by the lamination. Along with Theorem 3.10, this is enough to establish that the co-landing locus of any FDL is non-empty. The vertices of polygons in an FDL are all periodic or preperiodic, which is the same as rational. According to [3], all rational rays land at periodic or preperiodic point in the Julia set. Kiwi makes it clear with his characterization of what it means for a ray to land that the landing point of a ray is in its impression. Thus, his corollary 1.2 applies, and the impression of the ray is its landing point. ∎
@@ 7 Loops in the closure of the pullback tree?
This justifies an interest in the finer relation on the FDL on the same level of the pullback tree. The finer relation might be a bit messy in higher degree, so make things easier to draw, we add a notion, which hopefully adds information (see 8.4 and 8.5).
@@ 8 Future Directions of Research
    Consider a pullback tree. Let $P_{1}$ be the polynomial of a lamination in the base of the pullback tree. For any polynomial, $P_{2}$, such that $P_{1}$ is a primitive renormalization of $P_{2}$, there exists another lamination in the base of the pullback tree that is the lamination of $P_{2}$. However, for any polynomial, $P_{3}$, such that $P_{1}$ is a satellite renormalization of $P_{3}$, there is not another lamination in the base of the pullback tree that is the lamination of $P_{3}$.
    The latter proposition seems more provable since a satellite bifurcation of the canonical pullback lamination coexists with the canonical. (Two laminations coexist if their union is a lamination [6].) Thus, it seems that we can not force there to be a periodic leaf placed in the necessary place.
  3. 3.
    The following formula is comparable to the Riemann-Hurwitz relation found in [12].
    Definition 8.3.
    Conjecture 8.4.
    In any FDL other than the root, the sum of the degree of all the gaps minus the number of gaps equals $d-1$. Thus, for any FDL, the trapped criticality plus the free criticality is $d-1$.
    Conjecture 8.5.
  4. 4.
    Conjecture 8.6.

### v2

@@ 5 What are the limit sets of FDL?
Of course, we wish to classify these sequences by how nice their limits are. For that purpose we summarize the definitions of a few terms which we use sparsely. The most important term is q-lamination, which is defined in [1] as a lamination having an equivalent relation that is laminational with the added convention that a leaf is in the lamination iff it is in the boundary of the convex hull of an equivalence class. Laminational, is also defined in [1], and refers to an equivalence relation that meets a number of conditions that are each necessary for the lamination to be realized by a polynomial in the sense contemplated here. One such condition is that the equivalence classes are finite. These terms generally presuppose that the lamination is invariant. Unclean is defined in Thurston as having 3 leaves meet at a point. Unclean laminations are not q, but they might be proper, in which case their equivalence relations are laminational. When we say that a lamination is not proper we mean that it has infinite equivalence classes. To summarize, an unclean sibling lamination can be cleaned into a q-lamination iff it is proper.
@@ 6 Are all FDL realized by a polynomial?
Kiwi’s theorem does not show that all rays land in the way described by the lamination but instead that the impression of each ray intersects as described by the lamination. Along with the fact that all FDL have at least one child, this is enough to establish that the co-landing locus of any FDL is non-empty. The vertices of polygons in an FDL are all periodic or preperiodic, which is the same as rational. According to [3], all rational rays land at periodic or preperiodic point in the Julia set. Kiwi makes it clear with his characterization of what it means for a ray to land that the landing point of a ray is in its impression. Thus, his corollary 1.2 applies, and the impression of the ray is its landing point. ∎
@@ 7 Loops in the closure of the pullback tree?
This justifies an interest in the finer relation on the FDL on the same level of the pullback tree. The finer relation might be a bit messy in higher degree, so make things easier to draw, we add a notion, which hopefully adds information.
@@
Definition 7.4.
By Lemma 2.13, the free criticality and the trapped criticality add to $d-1$.
Conjecture 7.5.